# Generating Fine Details of Entity Interactions

### v1

@@ Abstract
Images not only depict objects but also encapsulate rich interactions between them. However, generating faithful and high-fidelity images involving multiple entities interacting with each other, is a long-standing challenge. While pre-trained text-to-image models are trained on large-scale datasets to follow diverse text instructions, they struggle to generate accurate interactions, likely due to the scarcity of training data for uncommon object interactions. This paper introduces InterActing, an interaction-focused dataset with 1000 fine-grained prompts covering three key scenarios: (1) functional and action-based interactions, (2) compositional spatial relationships, and (3) multi-subject interactions. To address interaction generation challenges, we propose a decomposition-augmented refinement procedure. Our approach, DetailScribe, built on Stable Diffusion 3.5, leverages LLMs to decompose interactions into finer-grained concepts, uses a VLM to critique generated images, and applies targeted interventions within the diffusion process in refinement. Automatic and human evaluations show significantly improved image quality, demonstrating the potential of enhanced inference strategies. Our dataset and code are available at https://concepts-ai.com/p/detailscribe/ to facilitate future exploration of interaction-rich image generation.
@@ 1 Introduction
In particular, many existing benchmarks for text-to-image models have been focusing on single objects or simple spatial relations. To address this, we propose a new dataset specifically curated for fine-grained and interaction-rich text-to-image generation. The dataset includes examples of functional and action-related interactions (e.g., using tools and making physical contacts), compositional relationships (e.g., geometric and abstract layouts), and multi-subject interactions. We evaluate models using a combination of vision-language model assessments, automatic metrics, and human evaluation protocols for a more comprehensive and robust measure of generation quality. Fig. 1 shows common failures of off-the-shelf T2I models [4] such as physically-feasible interactions and layout errors.
To enhance fine-grained interaction generation in T2I models, we introduce DetailScribe, a generate-then-refine framework based on concept decomposition. DetailScribe is the first framework to combine multi-modal LLMs’ reasoning (concept decomposition) and recognition (image critique) ability to improve text-to-image generation. It is compatible with most T2I models, preserves their diversity, and requires no additional datasets or domain-specific knowledge. At a high level, DetailScribe has two steps: generating an initial image from a prompt, followed by iterative refinement based on VLM critiques. Crucially, to enhance the critique process, we first prompt a large language model (LLM) to decompose the user-input prompt into a structured hierarchy of entities, attributes, and their interactions. This breakdown provides a more detailed and organized description of the scene. Based on this decomposed description, a VLM generates critiques against the generated image by identifying specific elements that deviate from the prompt. Using this feedback, we add additional noise to the generated image and apply diffusion-denoising steps for a second iteration. By steering the refined instruction on the critical components of the original prompt, this process allows precise adjustments to align the output more closely with the intended scene while preserving correct details in the generated image from the previous step.
2 Related Works
Text-to-image diffusion models. Diffusion models have significantly advanced the field of image generation [9]. Recent innovations in these models have pushed the boundaries of quality and fidelity in text-to-image generations [24, 25, 10, 4, 15]. However, challenges remain in handling complex relationships and intricate compositional structures, particularly when generating images that involve multiple subjects.
Another approach to generating fine details of object interactions is by adding additional images at inference time, also known as customization [14, 13]. They use test-time adaptation to learn a new concept, such as a particular relationship between two entities, and apply it in novel contexts. Their approach is orthogonal to ours because we do not rely on any user-provided images besides the single text prompt.
Text-to-image benchmarks. There has been work on the evaluation of text-to-image generation. Most of the existing benchmarks for image generation have focused on automated metrics to assess image quality and alignment, employing datasets such as MS-COCO [20] and ImageNet [2]. Some commonly used automated metrics are Inception Score [26], Fréchet Inception Distance (FID) [8], and CLIPScore [7]. Some human preference [19] studies have been conducted by requesting users to rank and rate image generation quality [29, 16]. Recently, Leeet al. [17] identified 12 aspects and 62 scenarios for a holistic evaluation. Huang et al. [12] proposed T2I-CompBench as a comprehensive compositional text-to-image (T2I) generation benchmark, consisting of text prompts from color binding, shape binding, texture binding, spatial relationships, non-spatial relationships, and complex compositions. Despite these valuable contributions, most existing works focus on the generation of detailed attributes that are explicitly described in the prompt, but still struggle to accurately interpret and infer the implicit interactions and relations of objects that require common-sense knowledge. In this work, we craft a dataset, named InterActing, to evaluate such ability of T2I models, and propose a method to enhance the inference time reasoning of the T2I generator.
@@ 2 Related Works
Table 1: The InterActing dataset contains 1000 text-to-image prompts. We categorize them into subclass and count the occurrence.
Inference scaling and self-correction. With the advancing capabilities of LLMs, these models have shown potential for evaluating the performance of other models or even themselves. During training, LLMs are often used to provide feedback rewards that improve alignment, as demonstrated by Rafailov et al. [23]. Recently, self-correction mechanisms have also been applied during inference to enhance LLM performance, albeit at the cost of increased computation. For example, Zhang et al. [30] iteratively prompts GPT to produce improved abstractive summaries; Gao et al. [6] reduces hallucinations through iterative revisions; Dong et al. [3] applies self-correction for code generation; and AlphaProof identifies mistakes in solving math problems to guide the model toward correct solutions. Pan et al. [22] offers a comprehensive review of these techniques. Most recently, Ma et al. [21] introduces an inference-time scaling framework for diffusion models by searching for optimal noise inputs.
Recent advancements have incorporated LLMs to control the generation of diffusion models[31, 18, 5]. However, efforts to adapt these self-correction capabilities for vision-language models (VLMs) have been limited. Wu et al. [28] is the first to explore this approach. Our method distinguishes itself by providing more granular feedback and attempting to retain partial diffusion steps, making the self-correction process more flexible yet reliable.
@@ 3 The InterActing Dataset
We proposed a new dataset, InterActing consisting of 1000 interaction-focused text prompts from 3 scenarios covering three major types or real-world interactions: functional relationships and action-based interactions, compositional spatial relationships, and multi-subject interactions. In contrast to existing efforts on text-to-image benchmarking focusing on single object generation [28], combination of spatial and attribute relationships [12], and holistic aspects such as aesthetics and multi-linguality [17], InterActing focuses on generation tasks involving entity interaction with non-trivial details.
  • Functional relationships and action-based interactions, including tool using and actions that involve rich physical contacts.
  • Compositional spatial relationships, which are usually presented in the form of objects forming abstract layouts of geometric patterns.
  • Multi-subject interactions of more than one entity.
Table 1 showcases examples of actions, layouts, and interaction patterns in the dataset. We have the statistics of the entire dataset in Appendix D.
Figure 2: The overall pipeline of DetailScribe. DetailScribe takes as input a single natural language instruction. It first prompts a large language model (LLM) to generate a breakdown of the concepts in the image, which guides a vision-language model (VLM) to attend to different regions of a generated image and suggests fixes. It then adds noises back to the generated image and re-runs the diffusion process with the VLM-refined prompt to generate a faithful and high-fidelity image with rich entity interactions.
@@ 3.1 Evaluation Metrics
Due to the challenges in assessing whether an image aligns with the prompt’s description, we primarily rely on human evaluation, referred to as the human Likert scale. We further explored the use of VLMs and pre-trained metrics for automatic evaluation purposes (Automatic evaluation.). We note that these evaluations are inherently more noisy, so we compared their agreement with human preferences on sampled image pairs generated by all models. We then use the most aligned scores and VLM questions as our auto-evaluators for benchmarking on the entire InterActing.
Human evaluation.
In the human evaluation process, we asked annotators to rate images on a Likert scale. For each prompt, annotators were presented with images generated by all models, which were randomly shuffled to mitigate order bias. Annotators were instructed to assign a score between 1 and 5 based on image-text alignment, following these guidelines:
  • 1: The image is completely irrelevant to the prompt.
  • 3: The image includes most relevant objects, but some elements implied by the prompt are missing (e.g., tools required for specific actions).
  • 4: The image captures most relevant objects and infers additional ones successfully, but there are minor issues with object relationships.
@@
We also leverage the reasoning abilities of VLMs for evaluation. The VLM evaluation process uses a prompt that includes the instruction, an image generated by a model, and the rating guidelines. The VLM evaluator then outputs a score on a scale of 1 to 5. Besides VLM, we adopted CLIPScore [7] and ImageReward [29] as pre-tained text-image matching metrics to assess the image-text alignment of the generation. We further include BLIP-VQA proposed in [12] to capture fine-grained text-image correspondences.
@@ 4 DetailScribe
In this section, we introduce our refinement-augmented generation framework (DetailScribe) for text-to-image generation modeling fine-grained entity interactions. Illustrated in Fig. 2, DetailScribe operates in three stages: 1) given an input natural language prompt, a large language model hierarchically decomposes it into detailed sub-concepts (Section 4.1); 2) an initial image is generated from the prompt using a text-to-image model, followed by a vision-language model critique conditioned on both the decomposed sub-concepts and the generated image (Section 4.2); 3) based on the critique, the prompt is refined and a re-denoising process corrects errors, yielding a more faithful and realistic generated image (Section 4.3).
Overall, our framework leverages a diffusion-based text-to-image model as the basic model for generating and refining images. Augmenting this basic model, DetailScribe is designed to handle highly variable user inputs that describe complex interactions between entities.
Figure 3: VLM-based critique and prompt refinement. Given the LLM-generated concept decomposition and an image generated using the user input, a vision-language model generates a critique of errors in the image, suggests corrections, and finally refines the prompt. This prompt will be used in a second-round diffusion process to refine the image.
@@ 4.1 Prompt Completion by Concept Decomposition
At the core of our pipeline, this module refines a user-provided natural language instruction by generating a more detailed, structured version of it. Specifically, we adopt the concept of visual abstraction schema proposed by Hsu et al. [11], which offers a concise structured scene representation in the form of a directed acyclic graph (DAG) of the original text input. Illustrated in Fig. 2, in a schema, each node represents a subcomponent of the higher-level concept, and dependencies between components are captured as edges within the graph. For example, the instruction “a hedgehog is rolling dough with a rolling pin” can be decomposed into distinct entities (e.g., the hedgehog, the dough, and the rolling pin) and their interactions (e.g., the hedgehog holding the rolling pin, and the rolling pin contacting the dough). Optionally, background elements like tables and windows can also be included.
This relational and hierarchical representation is highly flexible in representing diverse visual scene descriptions. Empirically, we find that explicitly performing this decomposition enhances the downstream vision-language models’ ability to capture fine-grained entity interactions, as it naturally provides a “checklist” for identifying errors and refining the prompt for more accurate visual generation. Similar to the findings from Hsu et al. [11], empirically, we found that large language models are proficient and reliable in generating such hierarchical and relational concept decompositions. Across all examples shown in this paper, we prompt with LLM with only a single in-context learning example (cooking). We provide details about our text prompts in the supplementary material.
4.2 Vision-Language Model-Based Critique and Prompt Refinement
At a high level, this module leverages a vision-language model (VLM) to refine the generated content iteratively. Initially, an image is generated using a base text-to-image model conditioned on the original user input. Next, we utilize a VLM (GPT-4o in our experiments) to critique the generated image. The VLM is prompted to review each element in the decomposed visual schema from Section 4.1, checking whether the generated image accurately reflects the specified entities and their interactions.
After the VLM identifies discrepancies, it proposes edits to the original prompt—typically by inserting a few focused phrases to enhance clarity or detail in specific parts of the sentence. This refined prompt is then used in a subsequent iteration to improve the generated image. As a concrete example, illustrated in Fig. 3, we prompt the VLM first to generate an itemized list of critiques, pointing to specific parts of the image, by referring to the LLM-generated concept decomposition. Next, based on the generated critiques, it suggests a new list of corrections. Finally, it generates a prompt by revising the user input, typically introducing additional details. In our implementation, these three steps are merged into a single prompt.
@@ 4.3 Refinement by Diffusion Re-denoising
In this subsection, we detail our framework for refining generated images based on vision-language model feedback using a re-denoising process. Our framework leverages the fact that diffusion models use progressive generation processes that iteratively add details to an image through a reverse diffusion process. Thus, to correct specific parts of an already-generated image, we can introduce controlled noise to the image and rerun the diffusion process with the updated prompt. This approach preserves the integrity of the existing content while selectively refining the areas that need adjustment.
Recall that a diffusion model is composed of two procedures: forward diffusion and reverse diffusion. In the forward diffusion phase, a clean image $I_{0}$ is gradually corrupted by adding Gaussian noise across a sequence of time steps $t=1,2,\cdots,T$. This is defined as:
$I_{t}=\sqrt{\bar{\alpha}_{t}}I_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathcal{N}(0,1),$
where $\bar{\alpha}_{t}$ controls the noise level at each time step, and $\mathcal{N}(0,1)$ represents standard Gaussian noise.
In the reverse diffusion phase, the process is inverted to recover the original image. Starting from a highly noisy sample $I_{T}$ sampled from Gaussian noise, the model iteratively denoises it based on a noise prediction model $\epsilon(I_{t},t)$. By conditioning this noise prediction model also on the text prompts, the model can generate images aligned with specific descriptions.
To incorporate the feedback from the VLM, we perform a partial re-denoising process to correct specific errors in the generated image. Instead of regenerating the entire image, we add controlled noise to the existing image such that it matches the noise level of a particular diffusion step $t^{\prime}$. This re-introduced noise perturbs the image just enough to allow for modifications while preserving its core structure. After adding noise, we rerun the reverse diffusion process using the refined prompt provided by the VLM, thereby selectively correcting details without losing the overall content. This framework effectively fine-tunes the image, ensuring it better aligns with the user’s original intent while maintaining coherence in the visual output.
Fig. 2 illustrates the refined image after applying our partial re-denoising process where we set $t^{\prime}=T-2$ for the Stable Diffusion 3.5 model. When $t^{\prime}$ is set to a very large value (e.g., $T$, corresponding to complete noise), although we can correct the identified errors, the entire image is effectively regenerated. This often introduces new errors. By contrast, if $t^{\prime}$ is set too low (i.e., when the image is already nearly clean), the updates from the refined prompt have minimal impact, making it difficult to incorporate necessary corrections. In general, for Stable Diffusion models, we found that setting $t^{\prime}$ close to $T-2$ gives the best performance. We ablate the choice of $t^{\prime}$ in detail in the experiment section.

### v2

@@ Abstract
Recent text-to-image models excel at generating high-quality object-centric images from instructions. However, images should also encapsulate rich interactions between objects, where existing models often fall short, likely due to limited training data and benchmarks for rare interactions. This paper explores a novel application of Multimodal Large Language Models (MLLMs) to benchmark and enhance the generation of interaction-rich images. We introduce InterActing, an interaction-focused dataset with 1000 LLM-generated fine-grained prompts for image generation covering (1) functional and action-based interactions, (2) multi-subject interactions, and (3) compositional spatial relationships. To address interaction-rich generation challenges, we propose a decomposition-augmented refinement procedure. Our approach, DetailScribe, leverages LLMs to decompose interactions into finer-grained concepts, uses an MLLM to critique generated images, and applies targeted refinements with a partial diffusion denoising process. Automatic and human evaluations show significantly improved image quality, demonstrating the potential of enhanced inference strategies. Our dataset and code are available at https://detailscribe.github.io/.
Generating Fine Details of Entity Interactions
Xinyi Gu  and Jiayuan Mao Massachusetts Institute of Technology {gxy, jiayuanm}@mit.edu
@@ 1 Introduction
Table 1: The InterActing dataset contains 1000 text-to-image prompts by subclass and occurrence.
In particular, many existing benchmarks for text-to-image models have been focusing on single objects or simple spatial relations. To address this, we propose a new Large language models (LLMs) generated dataset specifically curated for fine-grained and interaction-rich text-to-image generation. The dataset includes examples of functional and action-related interactions (e.g., using tools and making physical contacts), multi-subject interactions, and compositional relationships (e.g., geometric and abstract layouts). We evaluate models using a combination of MLLM assessments, automatic metrics, and human evaluation protocols for a more comprehensive and robust measure of generation quality. Fig. 1 shows common failures of off-the-shelf T2I models (Esser et al., 2024) such as physically-feasible interactions and layout errors.
To enhance fine-grained interaction generation in T2I models, we introduce DetailScribe, a generate-then-refine framework based on concept decomposition. DetailScribe is the first framework to combine multi-modal LLMs’ reasoning (concept decomposition) and recognition (image critique) ability to improve text-to-image generation. It is compatible with most T2I models, preserves their diversity, and requires no additional datasets or domain-specific knowledge. At a high level, DetailScribe has two steps: generating an initial image from a prompt, followed by iterative refinement based on MLLM critiques.
2 Related Works
Figure 2: The overall pipeline of DetailScribe. DetailScribe takes as input a single natural language instruction. It first prompts a large language model (LLM) to generate a breakdown of the concepts in the image, which guides a multimodal LLM (MLLM) to attend to different regions of a generated image and suggests fixes. It then adds noises back to the generated image and re-runs the diffusion process with the MLLM-refined prompt.
Text-to-image diffusion models. Diffusion models have significantly advanced the field of image generation, especially image fidelity (Ho et al., 2020; Ramesh et al., 2022; Saharia et al., 2022; Ho et al., 2022; Esser et al., 2024; Imagen-Team-Google et al., 2024). However, challenges remain in handling complex relationships and intricate compositional structures. Another approach to improving fine details of object interactions is by adding additional images at inference time (i.e., customization Huang et al., 2023b, 2024). Yet their approach is orthogonal to ours because we do not rely on any user-provided images.
Text-to-image benchmarks. Most of the existing image-generation benchmarks focus on automated metrics such as Inception Score (Salimans et al., 2016), FID (Heusel et al., 2017), and CLIPScore (Hessel et al., 2022) to assess image quality and alignment on image datasets such as MS-COCO (Lin et al., 2014) and ImageNet (Deng et al., 2009). Some human preference (Liang et al., 2024) studies have been conducted by requesting users to rank and rate images (Xu et al., 2023; Kirstain et al., 2023). Recently, more comprehensive synthetic benchmark datasets have been proposed for holistic evaluation (Lee et al., 2023; Huang et al., 2023a). Nevertheless, most existing works focus on the generation of detailed attributes that are explicitly described in the prompt, while still struggling to accurately modeling interactions among objects. In this work, we craft InterActing to evaluate such ability of T2I models.
Inference scaling and self-correction. MLLMs have shown potential for evaluating the performance of other models or even themselves. During training, MLLMs are often used to provide feedback rewards that improve alignment Rafailov et al. (2024). During inference, self-correction mechanisms have also been applied for improved performance (Pan et al., 2023; Zhang et al., 2023a; Gao et al., 2023; Dong et al., 2024; DeepMind, 2024). Most recently,  Ma et al. (2025) introduces an inference-time scaling framework for diffusion models by searching for optimal noise inputs. Recent advancements have incorporated MLLMs as agents to control the generation of diffusion models (Zhang et al., 2023b; Lian et al., 2024; Feng et al., 2023; Wu et al., 2024; Cho et al., 2023; Wang et al., 2024). Most works still rely on explicit object grounding and struggle to identify the fine-grained sub-part interactions. Our method focuses on more granular feedback and attempting to retain partial diffusion steps for better performance.
@@ 3 The InterActing Dataset
We proposed a new dataset generated by LLM, InterActing consisting of 1000 interaction-focused text prompts from 3 scenarios covering three major types or real-world interactions: functional relationships and action-based interactions (600), including tool manipulation and actions that involve rich physical contact, multi-subject interactions (200), and compositional spatial relationships (200), which are usually presented in the form of objects forming abstract layouts of geometric patterns. In contrast to existing efforts on text-to-image benchmarking focusing on single object generation (Wu et al., 2024), combination of spatial and attribute relationships (Huang et al., 2023a), and holistic aspects such as aesthetics and multi-linguality (Lee et al., 2023), InterActing focuses on tasks involving entity interaction with non-trivial details. We include more examples and statistics of the entire dataset in Appendix E.
@@ 3.1 Evaluation Metrics
Due to the challenges in assessing whether an image aligns with the prompt’s description, we primarily rely on human evaluation, referred to as the human Likert scale. Human annotators were instructed to assign a score between 1 and 5 based on image-text alignment. Detailed evaluation guidelines are listed Appendix C.1. We further explored the use of MLLMs and pre-trained metrics for automatic evaluation purposes (automatic evaluation). We note that these evaluations are inherently noisier, so we compared their agreement with human preferences on sampled image pairs generated by all models. We then use the most aligned scores and MLLM questions as our auto-evaluators for benchmarking on the entire InterActing.
@@
We also leverage the reasoning abilities of MLLMs for evaluation. The MLLM evaluation process uses a prompt that includes the rating instruction, and the image generated by a model (Appendix C.2). The MLLM evaluator then outputs a score on a scale of 1 to 5. Besides MLLM, we adopted CLIPScore Hessel et al. (2022) and ImageReward Xu et al. (2023) as pre-tained text-image matching metrics to assess the image-text alignment of the generation. We further include BLIP-VQA proposed in Huang et al. (2023a) to capture fine-grained text-image correspondences.
Figure 3: MLLM-based critique and prompt refinement. Given the LLM-generated concept decomposition and an image generated using the user input, an MLLM generates a critique of errors in the image, suggests corrections, and finally refines the prompt. This prompt will be used in a second-round diffusion process.
@@ 4 DetailScribe
In this section, we introduce our refinement-augmented generation framework (DetailScribe) for text-to-image generation modeling fine-grained entity interactions. Illustrated in Fig. 2, DetailScribe operates in three stages: 1) given an input natural language prompt, a large language model hierarchically decomposes it into detailed sub-concepts (Section 4.1); 2) an initial image is generated from the prompt using a text-to-image model, followed by an MLLM critique conditioned on both the decomposed sub-concepts and the generated image (Section 4.2); 3) based on the critique, the prompt is refined and a re-denoising process corrects errors, yielding a more faithful and realistic generated image (Section 4.3).
Overall, our framework leverages a diffusion-based text-to-image model as the base model for generating and refining images. Augmenting this base model, DetailScribe is designed to handle highly variable user inputs that describe complex interactions between entities.
@@ 4.1 Prompt Completion by Concept Decomposition
At the core of our pipeline, this module refines a user-provided natural language instruction by generating a more detailed, structured version of it. Specifically, we adopt the concept of visual abstraction schema proposed by Hsu et al. Hsu et al. (2024), which represents scene structures with a directed acyclic graph (DAG). Illustrated in Fig. 2, in a schema, each node represents a subcomponent of the higher-level concept, and dependencies between components are captured as edges. For example, the instruction “a hedgehog is rolling dough with a rolling pin” can be decomposed into distinct entities (e.g., the hedgehog, the dough, and the rolling pin) and their interactions (e.g., the hedgehog holding the rolling pin, and the rolling pin contacting the dough). Optionally, background elements like tables and windows can be included.
This relational and hierarchical representation is highly flexible, and we find that explicitly performing this decomposition enhances the downstream MLLMs’ ability to capture fine-grained entity interactions, as it naturally provides a “checklist” for identifying errors and refining the prompt.
Across all examples shown in this paper, we prompt LLM with only a single in-context learning example (cooking). We provide details about our text prompts in Appendix A.
Figure 4: Images generated by DetailScribe and baselines on the InterActing dataset. From left to right: 1) Stable Diffusion; 2) SD + GPT Rewrite; 3) SD + GPT Refine; 4) SD by Alternative seed/initial noise in the Inference scaling experiment; 5) DALL·E 3; 6) Ours: DetailScribe. DetailScribe consistently provides effective corrections, which help generate images that closely follow the fine details in the prompts.
4.2 Multimodal LLM-Based Critique and Prompt Refinement
At a high level, this module leverages a multimodal LLM (MLLM) to refine the generated image iteratively. Initially, an image is generated using a base T2I model conditioned on the original user input. Next, we utilize an MLLM (GPT-4o in our experiments) to critique the generated image. The MLLM is prompted to review each element in the schema from Section 4.1, checking whether the generated image accurately reflects the specified entities and their interactions.
After the MLLM identifies discrepancies, it proposes edits to the original prompt—typically by inserting a few focused phrases to enhance clarity or detail in specific parts of the sentence. This refined prompt is then used in a subsequent iteration to improve the generated image. Fig. 3 illustrated a concrete example for this step.
@@ 4.3 Refinement by Diffusion Re-denoising
Based on the MLLM feedback, our framework refines the image based on the fact that diffusion models iteratively add details to an image through a reverse diffusion process. Thus, to correct specific parts of an already-generated image, we can introduce controlled noise to the image and rerun the diffusion process with the updated prompt. This approach preserves the integrity of the existing content while selectively refining certain areas.
Recall that a diffusion model is composed of two procedures: forward diffusion and reverse diffusion. In the forward diffusion phase, a clean image $I_{0}$ is gradually corrupted by adding Gaussian noise across a sequence of time steps $t=1,2,\cdots,T$: $I_{t}=\sqrt{\bar{\alpha}_{t}}I_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathcal{N}(0,1),$ where $\bar{\alpha}_{t}$ controls the noise level at each time step, and $\mathcal{N}(0,1)$ represents standard Gaussian noise.
The reverse diffusion process recovers the original image from a Gaussian noise. The model iteratively denoises it based on a noise prediction model $\epsilon(I_{t},t)$. By conditioning this noise prediction model also on the text prompts, the model can generate images aligned with specific descriptions.
To incorporate the feedback from the MLLM, we perform a partial re-denoising process. Instead of regenerating the entire image, we add controlled noise to the existing image such that it matches the noise level of a particular diffusion step $t^{\prime}$. This noise perturbs the image just enough to allow for modifications while preserving its core structure. Next, we rerun the reverse diffusion process using the refined prompt from MLLM to selectively correct details without losing the overall content. Fig. 2 shows an refined image with $t^{\prime}=T-2$. We ablate the choice of $t^{\prime}$ in Appendix 5.4.